Peter saw Lara Jean running on the school's track field. He accelerates from rest to 9.7 m/s within 23.8 s until he catches up with her. If Peter's mass is 74.02 kg, what is the average power (in hp) dissipated by Peter's body?


1
Expert's answer
2021-03-04T17:24:44-0500

The average acceleration of Peter is:


"a = \dfrac{v_2m/s-0m/s}{t} = \dfrac{v_2}{t}"

where "v_2 = 9.7m\/s" is the final speed, "0m\/s" is the initial spee, and "t = 23.8s" is the time.

Distance covered under this acceleration is:


"d = \dfrac{at^2}{2} = \dfrac{v_2t}{2}"

The average net force acted on Peter's body is (by second Newton's law):


"F = ma"

where "m = 74.02kg" is Peter's mass. The average work done by Peter's body is:


"A = Fd = ma\dfrac{v_2t}{2} = \dfrac{mv_2^2}{2}"

Thus, the average power is given as (by definition):


"P = \dfrac{A}{t} = \dfrac{mv^2_2}{2t}\\
P = \dfrac{74.02\cdot 9.7^2}{2\cdot 23.8} \approx 146.3W"

Answer. 146.3 W.